OpenStudy (anonymous):

The cylinder tube of a spray pump has a radius R , one end of which has n fine holes each of radius r . If the speed of flow of the liquid in the tube is v , the speed of ejection through the hole is

OpenStudy (anonymous):

This is an imcomplete

OpenStudy (anonymous):

question

OpenStudy (anonymous):

No this is a complete question

OpenStudy (anonymous):

theres no values? what are the answers

OpenStudy (anonymous):

is this the question? http://screencast.com/t/v0DOiF2SjhM

OpenStudy (anonymous):

V/n[R/r]^2

OpenStudy (irishboy123):

use continuity \(\Large V_{tube} = \pi R^2l\) \(\Large \dot V_{tube} = \pi R^2 \dot l = \pi R^2 v \) \(\Large = n \times \pi r^2 \tilde v\), ie due to continuity, flow through holes must be same \(\Large \tilde v = \dfrac{\pi R^2 v}{ n \pi r^2} = \dfrac{v}{ n}. \dfrac{ R^2 }{ r^2}\)